# Reciprocal Beyond-Diagonal Reconfigurable Intelligent Surface  (BD-RIS):  Scattering Matrix Design via Manifold Optimization

### v3

@@ II-B Problem Formulation
Given $\mathbf{\Theta}$ and $\mathbf{V}$, and assuming perfect channel state information, an achievable rate at $U_{k}$ is $\log_{2}(1+\gamma_{k})$, where the signal to interference-plus-noise ratio (SINR) is expressed as
@@
where $a\in\{\text{SC},\text{FC},\text{GC}\}$, as described in [26].
The optimization problem (P1) is inherently non-convex due to the objective function in (10a), and the strong coupling between the TX beamforming and the scattering matrix. Additionally, the constraints in (10b)-(10d) are generally non-convex. As previously mentioned, the constraints in (10c) and (10d) reflect the BD-RIS’s reciprocal and lossless properties, respectively, with the constraint in (10d) representing the Stiefel manifold condition.
Due to this coupling, a two-stage design is adopted to enhance the sum-rate. In this article, the first stage of the optimization problem is addressed, namely, the design of the scattering matrix. The subsequent stage, focusing on beamforming matrix design, will be presented in a follow-up work.
@@ III Stage 1: Scattering Matrix Design
We will solve (P1) in an iterative manner. To first design the scattering matrix, the impact of the BS beamforming is fixed by initializing it with a minimum mean square error (MMSE)-based beamformer. Accordingly, (2f) is rewritten as
@@
where $\mathbf{\Omega}=\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}=[\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{N}]\in\mathbb{C}^{R\times N}$ and $\mathbf{E}=\mathbf{H}_{\mathrm{RX}}\mathbf{\Omega}=[\mathbf{e}_{1},\dots,\mathbf{e}_{K}]^{T}\in\mathbb{C}^{K\times N}$ denotes the equivalent channel. As a consequence, the SINR at $U_{k}$ during the design of the scattering matrix is given as
@@
Making use of (12), the optimization problem is formulated as
@@ III-A Solution for the Group-Connected architecture
As the group-connected BD-RIS architecture is the most generalized version of the three architectures considered, it can provide a generic solution to (P2), which in turn can be rewritten as
@@
where $g=\{1,2,\dots,G\}$. For simplicity, the sum-rate over all users is denoted as
@@
where $\nu\in\mathbb{R}$ denotes a nonnegative weight and $\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ is used to obtain a quantitative measure of how far the matrix is from being symmetric.
@@
Having defined the optimization problem (P3), its solution can be obtained numerically using the Manopt toolbox [36]. However, to reduce the high computational complexity associated with Manopt, the corresponding CGA algorithm [37], tailored for this problem, is derived in the sequel. This algorithm incorporates a closed-form expression for the gradient of the objective function, as seen in (16a), enabling a more efficient implementation of the scattering matrix design.
To this extent, the group-wise definition of the system model is considered, namely
where $g\in\{1,2,\dots,G\}$.
As such, the group-wise SINR  [^3] This notation considers only the contributions of $\mathbf{\Theta}_{g}$, with the contributions from all other groups set to zero. equation can be written as
$\gamma_{k}^{(g)}=\frac{\left|\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}},$ (18)
where $\mathbf{v}_{l}\in\mathbb{C}^{N\times 1}$ denotes the beamforming vector, with $l\in\{i,k\}$.
@@
The vector $\mathbf{h}_{k}^{(g)T}=\mathbf{H}_{RX\left[k,R_{G}(g-1)+1:gR_{G}\right]}$ and the matrix $\mathbf{W}^{(g)}=\mathbf{H}_{TX\left[R_{G}(g-1)+1:gR_{G},1:N\right]}$ represent the channel components associated with the $g$-th group. Specifically, $\mathbf{h}_{k}^{(g)T}\in\mathbb{C}^{1\times R_{G}}$ denotes the BD-RIS-$U_{k}$ channel vector, and $\mathbf{W}^{(g)}\in\mathbb{C}^{R_{G}\times N}$ denotes the BS-BD-RIS channel matrix corresponding to the group size $R_{G}$, respectively. Within the $g$-th group dimension, the scattering matrix $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$ models the group-specific scattering behavior. Accordingly, the group-wise equivalent channel is defined as $\mathbf{e}_{k}^{(g)}=\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}$, where $\mathbf{e}_{k}^{(g)}\in\mathbb{C}^{1\times N}$, enabling the SINR to be rewritten as
$\gamma_{k}^{(g)}=\frac{\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}}.$ (19)
@@
The gradient of the objective function (16a) with respect to each group $\mathbf{\Theta}_{g}$ is computed as
$\!\!\!\nabla_{\mathbf{\Theta}_{g}}\!\eta$ $=\nabla_{\mathbf{\Theta}_{g}}\!\!\left(\!\sum_{k}\log_{2}(1+\gamma_{k}^{(g)})-\nu\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}\!\right)$ (20)
$=\frac{1}{\ln(2)}\sum_{k}\!\frac{1}{1+\gamma_{k}^{(g)}}\nabla_{\mathbf{\Theta}_{g}}\gamma_{k}^{(g)}\!-\nu\nabla_{\mathbf{\Theta}_{g}}\!\!\left\|\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}\!,$
@@
The gradient for half of the terms in (21) can be effectively found in [38], namely
@@
For simplicity, the numerator and denominator of $\gamma_{k}$ in (19) are denoted as $\chi_{k}$ and $\psi_{k}$, respectively, $i.e.$,
$\gamma_{k}^{(g)}=\frac{\chi_{k}}{\psi_{k}}=\frac{\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}},$ (26)
such that the gradient of $\gamma_{k}$ can be expressed as
$\nabla_{\mathbf{\Theta}_{g}}\gamma_{k}^{(g)}=\frac{\psi_{k}\nabla_{\mathbf{\Theta}_{g}}\chi_{k}-\chi_{k}\nabla_{\mathbf{\Theta}_{g}}\psi_{k}}{\psi_{k}^{2}},$ (27)
where the gradients of the numerator and denominator are computed as follows:
$\nabla_{\mathbf{\Theta}_{g}}\chi_{k}=\nabla_{\mathbf{\Theta}_{g}}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}=2\,\Bigg((\mathbf{e}_{k}^{(g)}\mathbf{v}_{k})^{\star}\frac{\partial\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}}{\partial\mathbf{\Theta}_{g}}\Bigg)^{\star},$ (28)
$\nabla_{\mathbf{\Theta}_{g}}\psi_{k}$ $=\nabla_{\mathbf{\Theta}_{g}}\!\left(\sum_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}\right)$ (29)
$=2\,\Bigg(\sum_{i\neq k}(\mathbf{e}_{k}^{(g)}\mathbf{v}_{i})^{\star}\frac{\partial\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}}{\partial\mathbf{\Theta}_{g}}\Bigg)^{\star},$
@@
Finally, the partial derivatives seen in (28) and (29) are computed as
$\hskip-5.38193pt\frac{\partial\mathbf{e}_{k}^{(g)}\mathbf{v}_{l}}{\partial\mathbf{\Theta}_{g}}$ $=\frac{\partial}{\partial\mathbf{\Theta}_{g}}\!\left(\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{l}\right)\!=\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{l})^{T}\!.$ (30)
Accordingly, the gradient of the objective function in (20) is reformulated as shown in (31). For convenience, (32) presents the gradient of the objective function assuming a simple power allocation matrix used for beamforming [^4] This assumption can only be made for fully-loaded systems ($K$ = $N$)., $i.e.$, $\mathbf{V}=\mathrm{diag}(\sqrt{v}_{1},\dots,\sqrt{v}_{K})$, where $v_{k}\in\mathbb{R},\forall k$. Both gradients are provided at the top of the following page.
$\nabla_{\mathbf{\Theta}_{g}}\eta=\!\sum_{k}\frac{2\Big(\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}\!\!+\!N_{0}\Big)\!\Big(\!(\mathbf{e}_{k}^{(g)}\mathbf{v}_{k})^{\star}\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{k})^{T}\Big)^{\star}\!\!-2\!\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}\!\!\sum\limits_{i\neq k}\!\Big(\!(\mathbf{e}_{k}^{(g)}\mathbf{v}_{i})^{\star}\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{i})^{T}\Big)^{\star}}{\ln(2)(1\!+\!\gamma_{k}^{(g)})\Big(\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}\!\!+N_{0}\Big)^{2}}\!-4\nu(\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T})$ (31)
$\nabla_{\mathbf{\Theta}_{g}}\eta=\sum_{k}\frac{2v_{k}\Big(v_{k}\sum\limits_{i\neq k}\left|e_{k,i}^{(g)}\right|^{2}\!\!+N_{0}\Big)\Big(e_{k,k}^{(g)\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{k}^{(g)T}\Big)^{\star}\!\!-2v_{k}^{2}\left|e_{k,k}^{(g)}\right|^{2}\!\sum\limits_{i\neq k}\Big(e_{k,i}^{(g)\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{i}^{(g)T}\Big)^{\star}}{\ln(2)(1\!+\!\gamma_{k}^{(g)})\Big(v_{k}\sum\limits_{i\neq k}\left|e_{k,i}^{(g)}\right|^{2}\!\!+N_{0}\Big)^{2}}\!-4\nu(\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T})$ (32)
@@
The complex Stiefel manifold is defined as
$\mathrm{St}(n,p;\mathbb{C})=\left\{\mathbf{X}\in\mathbb{C}^{n\times p}\,:\,\mathbf{X}^{H}\mathbf{X}=\mathbf{I}_{p}\right\}.$ (33)
For square blocks ($p=n=R_{G}$), the Stiefel manifold $\mathrm{St}(R_{G},R_{G})$ is given by
$\mathrm{St}(R_{G},R_{G};\mathbb{C})=\{\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}:\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}}\}.$ (34)
@@
$\hskip-1.89444pt\boldsymbol{\Theta}=\mathrm{blkdiag}(\boldsymbol{\Theta}_{1},\boldsymbol{\Theta}_{2},\dots,\boldsymbol{\Theta}_{G}),\>\boldsymbol{\Theta}_{g}\in\mathrm{St}(R_{G},R_{G};\mathbb{C}),$ (35)
@@
$\mathrm{R}_{\mathbf{\Theta}}(\mathbf{\Theta},\boldsymbol{\Xi})=\mathcal{Q}(\mathbf{\Theta}+\alpha\boldsymbol{\Xi}),$ (36)
@@
$\mathbf{Y}_{g}=\mathbf{\Theta}_{g}+\alpha\boldsymbol{\Xi}_{g},\>[\mathcal{Q}_{g},\mathcal{R}_{g}]=\mathcal{QR}(\mathbf{Y}_{g}),\>\mathbf{\Theta}_{g}^{\text{new}}=\mathcal{Q}_{g}.$ (37)
@@
After the optimization, $i.e.$, following the sequence of retraction and tangent projection operations, the method from [27] is applied to enforce the symmetry and unitary constraints, since adding a penalty serves only as a workaround approach.
For the reader’s convenience, the projection method is summarized as follows. First, a symmetric matrix projection ensures that $\mathbf{A}=\mathbf{A}^{T}$, by averaging the solution with its transpose, $i.e.$,
$\Pi_{S_{sym}}(\mathbf{A})=\frac{1}{2}(\mathbf{A}+\mathbf{A}^{T}).\vskip-2.15277pt$ (40)
Second, to enforce the unitary constraint, the matrix is projected onto the set of unitary matrices using the singular value decomposition (SVD). Specifically, if $\mathbf{A}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{H}$, the unitary projection is defined as
$\Pi_{S_{uni}}(\mathbf{A})=\mathbf{U}\mathbf{V}^{H}.\vskip-2.15277pt$ (41)
Lastly, when both symmetry and unitary constraints must be enforced simultaneously, the procedure is as follows. If $\mathbf{B}$ is defined as the result after the symmetric projection of $\mathbf{A}$, $i.e.$, $\mathbf{B}=\Pi_{S_{sym}}(\mathbf{A})$, then given that the SVD of $\mathbf{B}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{H}$, the unitary symmetric projection of $\mathbf{A}$ is defined as
$\Pi_{S_{unisym}}(\mathbf{A})=\mathbf{U}\mathbf{V}^{H},$ (42)
where $\mathbf{U}$ and $\mathbf{V}^{H}$ denote the left and right singular vector matrices, respectively.
Algorithm 1 Proposed CGA for BD-RIS Optimization

### v4

@@ II-B Problem Formulation
Given $\mathbf{\Theta}$ and $\mathbf{V}$, and assuming perfect CSI  [^7] Although idealized, the assumption of perfect CSI is widely adopted in the RIS literature [14, 21, 34], as it enables evaluation of the “best-case” performance of the proposed algorithm, while imperfect CSI can only degrade performance. Moreover, in block-static channel conditions where the channel remains constant over the data transmission interval, the channel coefficients can, in principle, be estimated with arbitrarily high accuracy given sufficient training resources. In such settings, performance analysis under perfect CSI is consistent with the standard block-fading framework, where coding is performed over long coherence intervals. In contrast, in rapidly varying fading environments, channel estimation errors become unavoidable and must be explicitly incorporated into the system design. Extensions to robust formulations under imperfect CSI are feasible and have been extensively studied in the literature (see, e.g., [38, 39]). However, such developments fall beyond the scope of the present work., an achievable rate at the $k$-th user is $\log_{2}(1+\gamma_{k})$, where the signal to interference-plus-noise ratio (SINR) [^8] For the extension to the general MU-MIMO case, the SINR in (11) would need to be revisited to incorporate receive combining at the user side. is expressed as
@@
where the constraint (12b) corresponds to the total transmit power limitation at the BS, ensuring that the Frobenius norm of the beamforming matrix $\mathbf{V}$ does not exceed the maximum allowable transmit power $P_{\mathrm{max}}$, $a\in\{\text{SC},\text{FC},\text{GC}\}$ denotes the connectivity level of the BD-RIS, and $\mathcal{S}_{a_{1}}$ and $\mathcal{S}_{a_{2}}$ represent the sets of scattering matrices that satisfy the reciprocal and lossless constraints, respectively.
The optimization problem (P1) is inherently non-convex due to the objective function in (12a), and the strong coupling between the TX beamforming and the scattering matrix. Additionally, the constraints in (12b)-(12d) are generally non-convex. In particular, the constraint in (12d) represents a Stiefel manifold condition [30], since it enforces orthonormality of the columns of the BD-RIS scattering matrix.
Due to this coupling, a two-stage design is adopted to enhance the sum-rate. In this article, the first stage of the optimization problem is addressed, namely, the design of the scattering matrix. The second stage, concerning the optimization of the transmit beamforming matrices, becomes particularly relevant in more general MU-MIMO settings where active beamforming plays a larger role in the achievable performance. Further discussion on this aspect, as well as on the extension to the general MU-MIMO case, is provided in the next subsection.
II-C A Note on the General MU-MIMO Case
Before we proceed, let us clarify that the method proposed in this article can be directly applied to MU-MIMO systems with relatively few modifications described below.
Consider a MU-cell-free massive MIMO (CF-mMIMO) system with $L$ access points (APs), each equipped with $N_{a}$ TX antennas, such that the total number of transmit antennas is $N_{t}=L\times N_{a}$, and $K$ users equipped with with $M$ RX antennas, aided by a BD-RIS with $R$ REs. The corresponding channel linking the $l$-th AP to the $k$-th user can be defined as
$\mathbf{E}_{l,k}=\mathbf{H}_{\mathrm{RX},k}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX},l}\in\mathbb{C}^{M\times N_{a}},$ (13)
where $\mathbf{H}_{\mathrm{RX},k}$ and $\mathbf{H}_{\mathrm{TX},l}$ denote the BD-RIS-to-user and AP-to-BD-RIS channel matrices, respectively.
Defining the equivalent channel to the $k$-th user as the concatenation of the channel matrices from all APs to the $k$-th user, we write $\mathbf{E}_{k}\triangleq[\mathbf{E}_{1,k},\mathbf{E}_{2,k},\dots,\mathbf{E}_{L,k}]\in\mathbb{C}^{M\times N_{t}}$. Accordingly, the complex baseband received signal at the $k$-th user, which generalizes (1), is given by
$\mathbf{r}_{k}=\overbrace{\,\mathbf{E}_{k}\mathbf{V}_{k}\mathbf{s}_{k}}^{\text{Intended signal}}\;+\!\!\!\!\!\overbrace{\!\!\!\!\!\sum_{k^{\prime}\in\mathcal{K}\backslash\{k\}}\!\!\!\!\!\!\mathbf{E}_{k}\mathbf{V}_{k^{\prime}}\mathbf{s}_{k^{\prime}}}^{\text{Downlink inter-user interference}}\!\!+\mathbf{n}_{k}\in\mathbb{C}^{M\times 1},$ (14)
where $\mathbf{n}_{k}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$.
For a given rate $\log_{2}\text{det}(\mathbf{I}_{M}+\mathbf{\Gamma}_{k})$, the SINR at the $k$-th user, $\mathbf{\Gamma}_{k}$, generalizes (11) and is expressed as
$\mathbf{\Gamma}_{k}=\mathbf{V}_{k}^{\mathrm{H}}\mathbf{E}_{k}^{\mathrm{H}}\mathbf{\Psi}_{k}^{-1}\mathbf{E}_{k}\mathbf{V}_{k}\in\mathbb{C}^{M\times M},$ (15a)
where
[…]
is the interference-plus-noise covariance matrix, and $\mathbf{V}_{k}$ denotes the beamforming matrix for the $k$-th user.
From the above, it is evident that the proposed framework applies directly to general MIMO scenarios, as previously mentioned. We emphasize, however, such a setting would obscure the isolated impact of the BD-RIS scattering matrix design. In addition, a comprehensive treatment of the MIMO case would not be complete without considering also the design of corresponding beamformers. For these reasons, we leave the joint design of TX beamforming and BD-RIS scattering matrix for future work, and continue hereafter addressing the MU-MISO case, for simplicity, but without loss of generality.
@@ III Stage 1: Scattering Matrix Design
We will solve (P1) in an iterative manner. To first design the scattering matrix, the impact of the BS beamforming is fixed by initializing it with a minimum mean square error (MMSE)-based beamformer [^9] Perfect CSI is assumed as discussed previously. However, MMSE precoding has been shown to be robust under imperfect CSI [40], so this assumption does not represent a fundamental weakness of the considered framework.. Accordingly, (2f) is rewritten as
@@
where $\mathbf{\Omega}=\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}=[\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{N}]\in\mathbb{C}^{R\times N}$ and $\mathbf{E}=\mathbf{H}_{\mathrm{RX}}\mathbf{\Omega}=[\mathbf{e}_{1},\dots,\mathbf{e}_{K}]^{T}\in\mathbb{C}^{K\times N}$ denotes the equivalent channel.
As a consequence, the SINR at the $k$-th user during the design of the scattering matrix is given as
@@
Making use of (17), the optimization problem of concern can be formulated as
@@ III-A Solution for the Group-Connected Architecture
As the group-connected BD-RIS architecture is the most general version considered, it can provide a generic solution to (P2), which in turn can be rewritten as
@@
where $g=\{1,2,\dots,G\}$.
For simplicity, the sum-rate over all users is denoted as
@@
where $\nu\in\mathbb{R}$ denotes a nonnegative weight [^10] The penalty parameter $\nu$ is selected empirically based on preliminary simulations. While systematic hyperparameter optimization techniques could be employed (see, e.g., [41, 42, 43]), the focus of this work is on the scattering matrix design framework itself rather than on optimal regularization tuning. Moreover, in the following sections, the proposed method is shown to outperform existing SotA approaches even with empirically chosen $\nu$, and further performance improvements may be achieved through adaptive or data-driven tuning strategies. and $\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ is used to obtain a quantitative measure of how far the matrix is from being symmetric.
Having defined the optimization problem (P3), its solution can be obtained numerically using the Manopt toolbox [44]. However, measures can be taken to reduce the complexity of the optimization problem. In particular, to address the challenges arising in part from the non-convexity of the sum-rate function, the objective can be transformed into a convex form using FP techniques [45, 46].
To this extent, we apply the Lagrangian Dual Transform (LDT) to the sum-rate objective in (P3), which leads to the equivalent formulation
$\bar{\eta}_{k}=\log_{2}(1+\tau_{k})-\frac{\tau_{k}}{\ln(2)}+\frac{1+\tau_{k}}{\ln(2)}\cdot\frac{\left|\mathbf{e}_{k}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}},\vskip-4.30554pt$ (22)
where the auxiliary variable $\tau_{k}\in\mathbb{C}$ denotes the Lagrange multiplier, s.t. $\tau_{k}=\gamma_{k}$.
The fractional term in the majorized rate expression $\bar{\eta}_{k}$ given in (22) can be further convexified by applying the Quadratic Transform (QT), resulting in the following convex reformulation of the sum-rate objective:
$\hat{\eta}_{k}=$ $\log_{2}(1+\tau_{k})-\frac{\tau_{k}}{\ln(2)}$ (23)
$+\frac{1+\tau_{k}}{\ln(2)}\Big[2\Re\{y_{k}^{\star}\mathbf{e}_{k}\mathbf{v}_{k}\}-\left|y_{k}\right|^{2}\big(\sum_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}\big)\Big],$
where $y_{k}\in\mathbb{C}$ is an auxiliary variable introduced by the QT, given as
$y_{k}=\frac{\mathbf{e}_{k}\mathbf{v}_{k}}{\sum\limits_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}}.$ (24)
In this manner, the optimization problem (P3) can be reformulated as
$\hskip-5.8125pt(\text{P3a}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\!\!\sum_{k}\hat{\eta}_{k}-\nu\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ (25a)
subject to $\!\!\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},$ (25b)
where, for convenience, the modified sum-rate objective function in (25a), including the penalty term, is denoted as $\breve{\eta}_{k}$.
@@
In addition, to alleviate the high computational burden associated with Manopt, a tailored CGA algorithm [47] is derived in the sequel. This algorithm employs a closed-form expression for the gradient of the objective function, as given in (25a), thereby enabling a more efficient implementation of the scattering matrix design.
@@
To derive a solution for the general group-connected case, we define the vector $\mathbf{h}_{k}^{(g)T}=\mathbf{H}_{RX\left[k,R_{G}(g-1)+1:gR_{G}\right]}$ and the matrix ${\mathbf{H}^{(g)}_{\mathrm{TX}}}=\mathbf{H}_{\mathrm{TX}\left[R_{G}(g-1)+1:gR_{G},1:N\right]}$, which represent the channel components associated with the $g$-th group. Specifically, $\mathbf{h}_{k}^{(g)T}\in\mathbb{C}^{1\times R_{G}}$ denotes the BD-RIS-to-user $k$ channel vector, and ${\mathbf{H}^{(g)}_{\mathrm{TX}}}\in\mathbb{C}^{R_{G}\times N}$ denotes the BS-to-BD-RIS channel matrix corresponding to the group size $R_{G}$, respectively. Within the $g$-th group dimension, the scattering matrix $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$ models the group-specific scattering behavior. Accordingly, the group-wise equivalent channel is defined as $\mathbf{e}_{k}^{(g)}=\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}{\mathbf{H}^{(g)}_{\mathrm{TX}}}$, where $\mathbf{e}_{k}^{(g)}\in\mathbb{C}^{1\times N}$.
@@
As such, the gradient of the objective function (25a) with respect to each group $\mathbf{\Theta}_{g}$ is computed as
$\nabla_{\mathbf{\Theta}_{g}}\!\breve{\eta}=\nabla_{\mathbf{\Theta}_{g}}\!\Big(\!\sum_{k}\hat{\eta}_{k}-\nu\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}\!\Big),\vskip-4.30554pt$ (26)
@@
The gradient for half of the terms in (27) can be effectively found in [48, eq. (248)], namely
@@
Omitting the constant terms of the equivalent sum-rate objective after the FP transformations allows for the gradient to be expressed as
$\nabla_{\mathbf{\Theta}_{g}}\hat{\eta}_{k}\!=\!\nabla_{\mathbf{\Theta}_{g}}\!\Big(2\Re\{y_{k}^{\star}\mathbf{e}_{k}\mathbf{v}_{k}\}\!-\!\left|y_{k}\right|^{2}\!\big(\sum_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}\!+\!N_{0}\big)\Big),\vskip-4.30554pt$ (32)
where the gradient with respect to both terms is given as
$\nabla_{\mathbf{\Theta}_{g}}2\Re\{y_{k}^{\star}\mathbf{e}_{k}\mathbf{v}_{k}\}=2\left(y_{k}^{\star}\frac{\partial\mathbf{e}_{k}\mathbf{v}_{k}}{\partial\mathbf{\Theta}_{g}}\right)^{\star},$ (33)
and
$\nabla_{\mathbf{\Theta}_{g}}\left|y_{k}\right|^{2}\big(\sum_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}\big)=2\left|y_{k}\right|^{2}\sum_{i}\left(\big(\mathbf{e}_{k}\mathbf{v}_{i}\big)^{\star}\frac{\partial\mathbf{e}_{k}\mathbf{v}_{i}}{\partial\mathbf{\Theta}_{g}}\right)^{\star},$ (34)
@@
Finally, the partial derivatives in (33) and (34) are given by
$\hskip-5.38193pt\frac{\partial\mathbf{e}_{k}\mathbf{v}_{l}}{\partial\mathbf{\Theta}_{g}}$ $=\frac{\partial}{\partial\mathbf{\Theta}_{g}}\!\left(\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{l}\right)\!=\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{l})^{T}\!,$ (35)
where $l\in\{i,k\}$.
Accordingly, the gradient of the objective function in (26) is reformulated as shown in (36). For convenience, (37) presents the gradient of the objective function assuming a simple power allocation matrix used for beamforming [^11] This assumption can only be made for fully-loaded systems ($K$ = $N$)., $i.e.$, $\mathbf{V}=\mathrm{diag}(\sqrt{v}_{1},\dots,\sqrt{v}_{K})$, where $v_{k}\in\mathbb{R},\forall k$. For convenience, both gradients are provided at the top of the following page.
${\nabla_{\mathbf{\Theta}_{g}}\breve{\eta}=\sum_{k}\frac{1+\tau_{k}}{\ln(2)}\Big[2\left(y_{k}^{\star}\mathbf{h}_{k}^{(g)}\big(\mathbf{W}^{(g)}\mathbf{v}_{k}\big)^{T}\right)^{\star}-2\left|y_{k}\right|^{2}\sum_{i}\left(\big(\mathbf{e}_{k}\mathbf{v}_{i}\big)^{\star}\mathbf{h}_{k}^{(g)}\big(\mathbf{W}^{(g)}\mathbf{v}_{i}\big)^{T}\right)^{\star}\Big]\!-\!4\left(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right).}$ (36)
${\nabla_{\mathbf{\Theta}_{g}}\breve{\eta}=\sum_{k}\frac{1+\tau_{k}}{\ln(2)}\Big[2v_{k}\big(y_{k}^{\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{k}^{(g)T}\big)^{\star}-2\left|y_{k}\right|^{2}\sum_{i}v_{i}^{2}\big(e_{k,i}^{\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{i}^{(g)T}\big)^{\star}\Big]-4\nu(\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T}).}$ (37)
@@
The complex Stiefel manifold corresponding to square blocks ($p=n=R_{G}$) is given by
$\mathrm{St}(R_{G},R_{G};\mathbb{C})=\{\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}:\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}}\}.\vskip-2.15277pt$ (38)
@@
$\hskip-1.89444pt\boldsymbol{\Theta}=\mathrm{blkdiag}(\boldsymbol{\Theta}_{1},\boldsymbol{\Theta}_{2},\dots,\boldsymbol{\Theta}_{G}),\>\boldsymbol{\Theta}_{g}\in\mathrm{St}(R_{G},R_{G};\mathbb{C}),\vskip-2.15277pt$ (39)
@@
$\mathrm{R}_{\mathbf{\Theta}}(\mathbf{\Theta},\boldsymbol{\Xi})=\mathcal{Q}(\mathbf{\Theta}+\alpha\boldsymbol{\Xi}),\vskip-2.15277pt$ (40)
@@
$\mathbf{Y}_{g}=\mathbf{\Theta}_{g}+\alpha\boldsymbol{\Xi}_{g},\>[\mathcal{Q}_{g},\mathcal{R}_{g}]=\mathcal{QR}(\mathbf{Y}_{g}),\>\mathbf{\Theta}_{g}^{\text{new}}=\mathcal{Q}_{g}.\vskip-2.15277pt$ (41)
@@
After the optimization, $i.e.$, following the sequence of retraction and tangent projection operations, the method from [22] is applied to enforce the symmetry and unitary constraints, since adding a penalty serves only as a workaround approach. The full procedure implemented in this article is outlined in Algorithm 1.
Algorithm 1 Proposed CGA for RBD-RIS Optimization